OpenStudy (anonymous):

please help me fanning and metal. :) Eliza started her savings account with $100. Each month she deposits $25 into her account. Write the function that describes the relationship between the amount Eliza deposits and the total amount in her account. Then, determine the average rate of change in the account from the 2nd month to the 10th month.

OpenStudy (texaschic101):

f(x) = 100 + 25x ===> your function 2nd month -- f(2) = 100 + 25(2) f(2) = 100 + 50 f(2) = 150 so when x is 2, y is 150.....(2,150) 10th month -- f(10) = 100 + 25(10) f(10) = 100 + 250 f(10) = 350 so when x = 10, y = 350....(10,350) (2,150)....x1 = 2 and y1 = 150 (10,350)...x2 = 10 and y2 = 350 (y2 - y1) / (x2 - x1) (350 - 150) / (10 - 2) 200/8 = 25 average rate of change is 25.....I think thats how to do it...because I think the average rate of change is the same as the slope

OpenStudy (anonymous):

thanks!!!! :)